Circular No. 8155

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVAE IN M31 Further to IAUC 7709, M. Fiaschi, F. Di Mille, and D. Tiveron, report

the discovery of two apparent novae in M31 (= NGC 224) from Hα CCD images. One new object is located at α = 0^{h}42^{m}58^{s}.34, δ = +41^{o}16′07′′.8 (equinox 2000.0), which is 158′′ east and 1′′ south of the center of M31; available Hα magnitudes: Jan. 19, [18; June 19.034 UT, 16.5; 20.035, 16.6. The second new object is located at α = 0^{h}42^{m}52^{s}.22, δ = +41^{o}13′54′′.2, which is 89′′ east and 134′′ south of the center of M31; available Hα magnitudes: Jan. 19, [18; June 19.034, 15.6; 20.035, 15.7. The reference stars for the Hα photometry are from Ciardullo et al. (1987, Ap.J. 318, 520). Generally, after maximum, the nova color index R–[Hα] = +2 and remains approximately constant; V –[Hα] tends to change from +2 to +3 with time.

XTE J1814–338 D. Steeghs, Harvard-Smithsonian Center for Astrophysics, reports on

optical spectroscopy of XTE J1814–338 (cf. IAUC 8144) on June 22.2 UT with the 6.5-m Magellan/Clay telescope (+ Boller & Chivens spectrograph) at Las Campanas: “Prominent hydrogen and helium emission lines were detected in the spectrum of the optical-counterpart candidate reported by Krauss et al. (IAUC 8154). The classic hallmarks of an interacting binary spectrum firmly establishes the optical counterpart of the accreting pulsar. The strongest spectral feature was double-peaked Hα emission with FWHM = 800 km/s and equivalent width 0.6 nm.”

SUPERNOVA 2003ge IN NGC 6122 T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003ge (cf. IAUC 8154), obtained by P. Berlind on June 24.23 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova about two weeks past maximum.

V4745 SAGITTARII G. Pojmanski, Warsaw University Astronomical Observatory, reports

that ‘All Sky Automated Survey’ images obtained at Las Campanas show that V4745 Sgr brightened from V = 11.1 on June 22.108 UT to V = 10.1 on June 23.417. This appears to be the third secondary brightness maximum since Apr. 18; a 2.1-mag increase in V occurred on May 31.

2003 June 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green